OpenStudy (anonymous):

solve each equation. if the equation is an identity or if it has no solution, write identity or no solution. 89 + x= 2 - 2x

OpenStudy (lgbasallote):

89 + x = 2 - 2x add 2x to both sides 89 + 2x +x = 2 subtract 89 from both sides 2x + x = 2 - 89 got it?

OpenStudy (anonymous):

would it be x= -29

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

thanks